A compound contains 46.7% N and 53.3% O. If the molecular mass of the compound is 60.0g/mol. What is the molecular formula?
1
Expert's answer
2021-10-24T01:28:10-0400

46.7% Nitrogen

53.3% Oxygen


Firstly divide by their respective molar mass

"\dfrac{46.7}{14}= 3.34" Nitrogen


"\dfrac{53.3}{16}= 3.33" Oxygen


Then Divide by the lowest result

"\dfrac{3.34}{3.33} \approx 1" Nitrogen


"\dfrac{3.33}{3.33} \approx 1" Oxygen



The empirical formula is therefore, "NO\ (N_1O_1)"


The molecular mass is 60.0 g/mol

"\therefore (NO)n = 60\\
(14+16)n = 60\\
(30)n = 60\\
n = 2"


"\therefore" The molecular formula is "N_2O_2"